OpenStudy (anonymous):

Solve the exponential equation algebraically. Approximate to the 5th decimal place: 2^(1+x) = 3

OpenStudy (anonymous):

take the log of both sides (1+x) log 2 = log 3 1+x = (log3)/log2 x= log3/log2 - 1 x=0.58496